OpenStudy (anonymous):

A 16μ F capacitor is joined in series with a 34μ F capacitor and a 12 Volts battery. What is the potential difference across the first capacitor?What is the charge on the first capacitor?How much energy is stored by the first capacitor?

OpenStudy (anonymous):

Find equivalence capacitance of the circuit, Ceq=(16*34)/(16+34)=10.88

OpenStudy (anonymous):

Ceq=(16*34)/(16+34)=10.88 micro faraday. Hence the Net charge flows in the circuit=Ceq*V=130.56 micro coulomb hence the charge on first capacitor=130.56 micro coulomb and potential difference across it=130.56/16=8.16Volt